### v1

# Epidemiology-informed Graph Neural Network for Heterogeneity-aware Epidemic Forecasting

Abstract
Among various spatio-temporal prediction tasks, epidemic forecasting plays a critical role in public health management. Recent studies have demonstrated the strong potential of spatio-temporal graph neural networks (STGNNs) in extracting heterogeneous spatio-temporal patterns for epidemic forecasting. However, most of these methods bear an over-simplified assumption that two locations (e.g., cities) with similar observed features in previous time steps will develop similar infection numbers in the future. In fact, for any epidemic disease, there exists strong heterogeneity of its intrinsic evolution mechanisms across geolocation and time, which can eventually lead to diverged infection numbers in two “similar” locations. However, such mechanistic heterogeneity is non-trivial to be captured due to the existence of numerous influencing factors like medical resource accessibility, virus mutations, mobility patterns, etc., most of which are spatio-temporal yet unreachable or even unobservable. To address this challenge, we propose a Heterogeneous Epidemic-Aware Transmission Graph Neural Network (HeatGNN), a novel epidemic forecasting framework. By binding the epidemiology mechanistic model into a GNN, HeatGNN learns epidemiology-informed location embeddings of different locations that reflect their own transmission mechanisms over time. With the time-varying mechanistic affinity graphs computed with the epidemiology-informed location embeddings, a heterogeneous transmission graph network is designed to encode the mechanistic heterogeneity among locations, providing additional predictive signals to facilitate accurate forecasting. Experiments on three benchmark datasets have revealed that HeatGNN outperforms various strong baselines. Moreover, our efficiency analysis verifies the real-world practicality of HeatGNN on datasets of different sizes.
Index Terms:
Spatio-temporal Graphs, Predictive Analytics, Epidemic Forecasting
I Introduction
The outbreak and spread of epidemics are massive disasters facing global human society, which can cause a substantial number of deaths or irreversible physiological damage to humans[1] as well as significant economic losses across various industries worldwide[2]. Recent global outbreaks, such as COVID-19, have placed unprecedented pressure on public health systems and decision-makers [3]. In response, there has been a growing interest in developing accurate and effective epidemic forecasting models[4, 5]. These models can support the implementation of public health interventions[6], optimize resource allocation[7], and enhance public awareness of early prevention and control strategies[8], thereby helping to prevent and control the onset and rapid spread of epidemics.
Figure 1: Spatio-temporal and mechanistic heterogeneity between different locations in epidemic transmission. (a) Epidemic transmission with different influencing factors. (b) Example of heterogeneous epidemic transmission between two locations. (c) The analyses of two heterogeneities on a real-world dataset (US-Regions).
In a nutshell, epidemic forecasting predicts the infection numbers at geographically connected regions in the future horizon based on existing records. Traditional epidemic models, like the Susceptible-Infected-Recovered (SIR) model, rely on mechanistic approaches and location networks to simulate multi-regional transmission for epidemic forecasting[9, 10, 11]. Despite advantageous explainability and applicability, these models exhibit limited capacity to capture real-world complexities. As such, with the rapid development of machine learning, methods represented by spatio-temporal graph neural networks (STGNNs) have been used to learn spatial and temporal dependencies in epidemic transmission simultaneously [5, 4, 12, 13, 14, 15, 16]. However, many STGNNs are purely data-driven methods that neglect the underlying epidemic transmission mechanisms, making them prone to overfitting historical data and being misled by spurious features. To address the limitations of both sides, epidemiological machine learning (EML) models [17, 18, 19] have been proposed by incorporating mechanistic laws into a machine learning backbone to enhance the model’s generalizability and robustness. This trend also extends to STGNNs, where the time- and location-sensitive parameters in epidemic models are predicted to support dynamic epidemic forecasting[20, 21].
Generally, a key to epidemic forecasting is heterogeneity discovery, such that similar spatio-temporal patterns among locations can be aligned to enhance predictions while dissimilar ones are filtered out. This also explains the popularity of STGNNs for this task owing to their capability of learning the semantic affinity of connected locations within the topology of the graph. Meanwhile, existing models commonly use an over-simplified assumption that such heterogeneity only exists within observed features, such as past infection number, population, and demographic statistics of a region. However, in reality, epidemic heterogeneity holds for not only spatio-temporal features, but also the underlying intrinsic epidemic transmission mechanism. In this paper, we refer to these two types of heterogeneity as spatio-temporal heterogeneity and mechanistic heterogeneity, respectively. For illustration, we provide an example in Figure 1. In a nutshell, while spatio-temporal heterogeneity compares observed features (i.e., what they are now), like infection numbers, the mechanistic heterogeneity highlights the evolutionary trajectories of each location (i.e., how they will change), such as infection rates. It is worth noting that the two heterogeneity types are not mutually substitutable but rather complementary. For instance, though the two locations in Figure 1(b) have similar patterns in their infection numbers until the 20th week, the numbers have moved toward different directions in the following two weeks because of the distinct dynamics in the two locations’ infection rates. In the real epidemic transmission process, the two heterogeneity signals can vary greatly, as depicted by the Pearson Correlation Coefficient (PCC) across locations from a sampled time step (Figure 1(c)). Hence, if only the spatio-temporal heterogeneity is considered in the STGNN, the model is prone to making inaccurate predictions in the long term.
To this end, bringing awareness of the mechanistic heterogeneity to the prediction model appears to be a beneficial move. However, learning such heterogeneity incurs non-trivial challenges. The first challenge lies in effectively understanding and mimicking the underlying intrinsic transmission mechanisms, which are jointly determined by a complex suite of factors. As shown in Figure 1(a), those influencing factors [22, 6, 23, 24, 25, 8, 26] include (but are not limited to) public health resources, policy interventions, climates, individual or group susceptibility, virus mutations, and so on, many of which are not linked to the epidemic records and are even unobservable. Consequently, this renders explicit and precise modeling of the intrinsic epidemic transmission mechanism infeasible. At the same time, the second challenge is the modeling of the similarity/dissimilarity across the epidemic transmission mechanism of distinct regions. The root cause is the lack of effective measures to represent and quantify the mechanistic heterogeneity across locations, preventing it from providing more predictive signals into epidemic forecasting.
To address these challenges, we propose an innovative framework for heterogeneous epidemic transmission forecasting modeling: Heterogeneous Epidemic-Aware Transmission Graph Neural Network (HeatGNN). This framework aims to improve epidemic forecasting performance by simultaneously learning spatio-temporal and mechanistic heterogeneity. Firstly, we introduce epidemiology-informed embedding learning to encapsulate the mechanistic patterns into each location’s representation. To bypass the intricacy and unavailability of all possible influencing factors, instead of simulating the explicit mechanistic process of epidemic transmission, we parameterize the time-varying SIR model for each location with its spatio-temporal graph embedding. By aligning the behaviors of the implicitly parameterized time-varying SIR model and the explicitly derived one, the epidemiology-informed embedding can be effectively optimized. Secondly, to quantify and take full advantage of the mechanistic heterogeneity across locations, we propose to learn the time-varying mechanistic affinity graphs. Each mechanistic affinity graph is constructed based on the pairwise similarity between two epidemiology-informed location embeddings, where a sparsification approach is in place to remove noisy connections and ensure efficiency. With a backbone STGNN, the spatio-temporal and mechanistic heterogeneities are captured by respectively modeling the geographical and mechanistic affinity graphs across time steps to facilitate heterogeneity-aware epidemic forecasting. In summary, our main contributions are as follows:
  • In the context of epidemic forecasting, we characterize two types of heterogeneity, namely, saptio-temporal heterogeneity and mechanistic heterogeneity. We point out that on top of the commonly used spatio-temporal heterogeneity, mechanistic heterogeneity is a largely overlooked yet crucial predictive signal for predicting future infection numbers.
  • We propose HeatGNN, a novel framework for heterogeneous epidemic forecasting that jointly discovers spatio-temporal and mechanistic heterogeneity. By incorporating a well-defined mechanistic model into the STGNN, HeatGNN learns epidemiology-informed location embedding and mechanistic affinity graphs to account for mechanistic heterogeneity across locations to greatly benefit the predictive performance.
  • Our framework achieves state-of-the-art performance across three real-world epidemic datasets. Our efficiency analysis demonstrates that HeatGNN’s inference time is highly scalable w.r.t. the number of locations involved in the task.
Figure 2: The schematic illustration of HeatGNN.
II Preliminary
II-A Problem Formalization
In epidemic transmission, we formulate the epidemic forecasting problem as a node-level regression task on graphs. There are $N$ locations of interest in an area; each location (e.g., a region, city, or state) is represented as a node. Given the spatio-temporal observations of epidemic transmission risk dynamics: $\boldsymbol{X}=\{\boldsymbol{x}_{1},\boldsymbol{x}_{2},\ldots,\boldsymbol{x}_{T}\}^{T}\in\mathbb{R}^{T\times N}$, where $T$ represents the length of time step (e.g., the infected population number or infection number for $T$ weeks). Each $\boldsymbol{x}_{t}=\{x_{t,1},x_{t,2},\ldots,x_{t,N}\}\in\mathbb{R}^{N\times 1}(t=1,2,\ldots,T)$ is an $N-$dimensional vector, with $x_{t,i}(i=1,2,\ldots,N)$ being the epidemic transmission risk at the time step $t$ in the location $i$. The epidemic is transmitted in different areas, let $\mathcal{G}=(\mathcal{V},\mathcal{E})$ be a geographical graph, where $\mathcal{V}$ is the nodes set comprising $|\mathcal{V}|=N$ nodes (e.g., regions, cities, or states). The $\mathcal{E}\subseteq\mathcal{V}\times\mathcal{V}$ is the geographic link between these nodes. The adjacency matrix $\mathcal{A}\in\mathbb{R}^{N\times N}$ of the geographical graph is formulated as $\mathcal{A}_{ij}=1$ if there is an edge $e_{ij}\in\mathcal{E}$, and $\mathcal{A}_{ij}=0$ otherwise.
The objective of this work is to accurately predict the infection number in each location. Given the historical data and the adjacency matrix of the geographical graph, we can develop a predictive model $f$, which can be formulated as follows:
$\boldsymbol{x}_{t+h}=f(\boldsymbol{X}_{t-w+1:t},\mathcal{A}),$ (1)
where $\boldsymbol{x}_{t+h}$ is an epidemic transmission risk in a future time step $t+h$ and $h$ represents the forecasting horizon (e.g., the number of time steps to be predicted ahead). $\boldsymbol{X}_{t-w+1:t}=\{\boldsymbol{x}_{t-w+1},\boldsymbol{x}_{t-w+2},\ldots,\boldsymbol{x}_{t}\}\in\mathbb{R}^{w\times N}$ is the historical epidemic transmission risk, $w$ represents the window size of historical input, and $N$ is number of locations.
III Methodologies
The proposed framework, as shown in Figure 2, consists of four modules: 1) Spatio-temporal Graph Learning (STGL) Module; 2) Epidemiology-informed Embedding Learning (EIEL) Module; 3) Transmission Graph (TG) Module; and 4) Epidemic Forecasting Decoder Module. In the following subsections, we describe each module in more detail.
III-A Spatio-temporal Graph Learning Module
The spread of an epidemic in a city is influenced by nearby cities with similar geographic features, as adjacent locations have similar weather and topographical conditions, leading to comparable epidemic transmission patterns. Conversely, distinct weather and topography can result in varying patterns of spread. These differences lead to the emergence of spatio-temporal heterogeneity in epidemic transmission. To capture spatio-temporal heterogeneity, we propose using the STGL module in epidemic forecasting. The STGL module utilizes historical data to learn dynamic time series variations and geographic interactions, enabling it to capture the spatio-temporal heterogeneity of epidemic transmission across locations. EpiGNN is a graph neural network-based epidemic forecasting model that considers local and global spatial effects of regions and captures spatio-temporal dependencies well in some datasets[12]. Therefore, we use the backbone of EpiGNN as our STGL module. Given the historical epidemic transmission risk $\boldsymbol{X}_{t-w:t}$ and the adjacency matrix $\mathcal{A}$ of geographic topology, the spatio-temporal (ST) embedding $\boldsymbol{P}_{t}\in\mathbb{R}^{N\times D_{1}}$ with $D_{1}$-dimensional containing spatio-temporal heterogeneity information in $N$ locations can be described as:
$\boldsymbol{P}_{t}=f_{\text{ST}}(\mathcal{A},\boldsymbol{X}_{t-w:t}),$ (2)
where $f_{\text{ST}}$ is the STGL module.
III-B Epidemiology-informed Embedding Learning Module
The STGL module only captures the spatio-temporal heterogeneity among different locations, while many previous approaches predict epidemic transmission directly based on the learned ST embedding. However, only capturing the spatio-temporal pattern of the infected population is noisy for epidemic forecasting and lacks the intrinsic epidemic transmission mechanism, which is prone to producing generalization errors in predictions[18]. Therefore, we use the epidemiology mechanistic model to enhance the process of learning location transmission dynamics to capture latent patterns and reflect well-defined disease dissemination mechanisms. Given a location $i$, we assign the time-varying SIR model as the epidemiology mechanistic model, which can be described as following ordinary differential equations (ODEs):
[…]
where $S_{i}$, $I_{i}$, and $R_{i}$ represent the number of susceptible, infected, and recovered populations in location $i$, respectively. $\beta_{i}$ and $\gamma_{i}$ are the time-dependent parameters, which mean the infection rate of susceptible populations and the recovery rate of infected populations in the location $i$. Here, we follow the assumption that the total population $N_{i}=S_{i}+I_{i}+R_{i}$ consists of susceptible, infected, and recovered populations.
Although the time-varying SIR model reflects the intrinsic epidemic transmission mechanism, it requires high computational costs and domain knowledge to predict its large-scale parameters simultaneously. In addition, the variables of the time-varying SIR model are heavily entangled, and it is difficult to focus only on predicting changes in the number of infected populations ($I$) in epidemic forecasting. These challenges motivate us to propose the EIEL module that uses node embeddings to parameterize the five time-varying SIR variables. The EIEL module uses five different multi-layer perceptrons (MLPs) to predict susceptible $S$, infected $I$, recovered $R$, infection rate $\beta$, and recovery rate $\gamma$ in the time-varying SIR model at each location and time step. Each MLP consists of a fully connected (FC) layer with $L$ layers and an output layer. Additionally, to enhance the embedding capacity of the EIEL module, we learn epidemiology-informed location embeddings based on the ST embedding $\boldsymbol{P}_{t}=\{\boldsymbol{p}_{t,1},\boldsymbol{p}_{t,2},\ldots,\boldsymbol{p}_{t,N}\}\in\mathbb{R}^{N\times D_{1}}$. Then, based on $i$-th location ST embedding $\boldsymbol{p}_{t,i}\in\mathbb{R}^{D_{1}}$, the epidemiology-informed location embedding $\boldsymbol{e}_{j,i}\in\mathbb{R}^{D_{2}}$ of $j$-th time-varying SIR variable in $i$-th location can be computed by the $j$-th FC layer with $L$ layers:
$\boldsymbol{e}_{j,i}=g_{j}\left(\boldsymbol{p}_{t,i};\{\boldsymbol{W}^{(l)}_{j},\boldsymbol{b}^{(l)}_{j},\sigma^{(l)}_{j}\}_{l=1}^{L}\right),$ (4)
where $\boldsymbol{W}_{j}^{(l)}$ is the weight matrix of the $l$-th hidden layer of the $j$-th FC layer, $\boldsymbol{b}_{j}^{(l)}$ is its bias vector, and $\sigma_{j}^{(l)}$ is the activation function. Through five FC layers, we obtain the epidemiology-informed location embeddings $\boldsymbol{h}_{t+h,i}=[\boldsymbol{e}_{1,i};\boldsymbol{e}_{2,i};\boldsymbol{e}_{3,i};\boldsymbol{e}_{4,i};\boldsymbol{e}_{5,i}]\in\mathcal{R}^{5D_{2}}$ for $i$-th location in time step $t+h$. To predict the dynamic epidemic transmission, the output of MLP is defined as follows:
$y_{j,i}=\sigma_{j}^{(2)}(\boldsymbol{W}_{j}^{(2)}\boldsymbol{e}_{j,i}+\boldsymbol{b}_{j}^{(2)}),$ (5)
where $y_{j,i}\in\{\hat{S}_{t+h,i},\hat{I}_{t+h,i},\hat{R}_{t+h,i},\hat{\beta}_{t+h,i},\hat{\gamma}_{t+h,i}\}$ one of five variables of the susceptible $S$, infected $I$, recovered $R$, infection rate $\beta$, and recovery rate $\gamma$ at location $i$ at time step $t+h$. $\boldsymbol{W}_{j}^{(2)}$, $\boldsymbol{b}_{j}^{(2)}$, and $\sigma_{j}^{(2)}$ are the weight matrix, bias vector, and activation function in the output layer of the $j$-th MLP, respectively. In Section III-E, we introduce the optimization of the EIEL module to ensure that the learned epidemiology-informed location embedding is consistent with the intrinsic epidemic transmission mechanism.
III-C Transmission Graph Module
III-C1 Mechanistic Affinity Graph
The EIEL captures the intrinsic epidemic transmission mechanism of each location separately but cannot learn the interactive relationships between them. To further capture the differences in intrinsic epidemic transmission mechanisms between locations, we propose a time-varying mechanistic affinity graph (MAG) that can efficiently characterize the mechanistic heterogeneity between locations. The MAG $\widetilde{\boldsymbol{M}}_{t}\in\mathbb{R}^{N\times N}$ is calculated by the cosine similarity function $\phi$ between the epidemiology-informed location embeddings of different locations. The mechanistic similarity between locations is defined as follows:
$m_{ij}=\phi(\boldsymbol{h}_{t,i},\boldsymbol{h}_{t,j}),$ (6)
where the mechanistic similarity $m_{ij}\in\boldsymbol{M}_{t}$ represent the mechanistic dependence between $i$-th location and $j$-th location based on epidemiology-informed location embeddings ($\boldsymbol{h}_{t,i}$ and $\boldsymbol{h}_{t,j}$) of two locations at time step $t$. $\boldsymbol{M}_{t}$ is a symmetric matrix. Then, we normalize the similarity between different locations:
$\widetilde{\boldsymbol{M}}_{t}=\boldsymbol{D}^{-\frac{1}{2}}\boldsymbol{M}_{t}\boldsymbol{D}^{-\frac{1}{2}},$ (7)
where $\boldsymbol{D}$ is degree matrix of $\boldsymbol{M}_{t}$. To focus on the key mechanistic dependence, reduce redundant information, and enhance the computational efficiency of GNNs, we design an indicator function to sparsify the MAG $\widetilde{\boldsymbol{M}}_{t}\in[0,1]$. This sparsification is controlled by a sparse threshold $\delta$, applied to each normalized similarity $\widetilde{m}_{ij}\in\widetilde{\boldsymbol{M}}_{t}$:
$\widetilde{m}_{ij}=\begin{cases}\widetilde{m}_{ij}&\text{ if }\widetilde{m}_{ij}\geq\delta\\ 0&\text{ otherwise. }\end{cases},$ (8)
III-C2 Heterogeneous Transmission Graph Network
To fully capture the complex mechanistic heterogeneity by utilizing the mechanistic dependence in the MAG, we propose the Heterogeneous Transmission Graph Network (HTGN) to encode the transmission embedding between locations. The HTGN integrates a MAG with the architecture of Spatio-Temporal Graph Convolutional Networks (STGCN) [27], which comprises two spatio-temporal convolutional blocks (ST-blocks). Given that the historical epidemic transmission risk $\boldsymbol{X}_{t-w+1:t}$ and a MAG $\widetilde{\boldsymbol{M}}_{t}$ in time step $t$, we learn the transmission embedding $\boldsymbol{T}_{t}\in\mathbb{R}^{N\times D_{3}}$ of mechanistic heterogeneity by ST-Blocks:
$\boldsymbol{T}_{t}=\text{ST-Blocks}(\widetilde{\boldsymbol{M}}_{t},\boldsymbol{X}_{t-w+1:t}),$ (9)
III-D Epidemic Forecasting Decoder Module
Finally, by simultaneously capturing the spatio-temporal heterogeneity and mechanistic heterogeneity across different locations, our framework applies an FC layer to predict future epidemic transmission risk, providing the epidemic forecasting result $\hat{\boldsymbol{Y}}_{t+h}$ in time step $t+h$:
$\hat{\boldsymbol{Y}}_{t+h}=\boldsymbol{W}[\boldsymbol{P}_{t};\boldsymbol{T}_{t}]+\boldsymbol{b},$ (10)
where $\boldsymbol{W}$ and $\boldsymbol{b}$ are the weight matrix and basis vector, respectively, $\boldsymbol{P}_{t}$ is the ST embedding learned from the STGL module and $\boldsymbol{T}_{t}$ is the transmission embedding learned from the G module.
III-E Optimization
To ensure the epidemiology-informed location embeddings are consistent with the mechanism of the intrinsic epidemic transmission, we use the idea from the PINN [28] to propose a physics loss to inform the learning process in the EIEL module. The physics loss incorporates the underlying physical laws of the intrinsic epidemic transmission mechanism into the EIEL module. Specifically, we aim to solve the time-varying SIR model using five MLPs by minimizing the multi-objective physics loss $\mathcal{L}_{p}$. Given the prediction of five variables ($\hat{S}_{t+h,i}$, $\hat{I}_{t+h,i}$, $\hat{R}_{t+h,i}$, $\hat{\beta}_{t+h,i}$, and $\hat{\gamma}_{t+h,i}$) of time-varying SIR model, the physics loss $\mathcal{L}_{p}$ to be described as follows:
$\begin{gathered}\mathcal{L}_{d}(\hat{I}_{t+h,i})=\frac{1}{N}\sum_{i=1}^{N}|\hat{I}_{t+h,i}-x_{t+h,i}|,\\ \mathcal{L}_{o}(\hat{S}_{t+h,i},\hat{I}_{t+h,i},\hat{\beta}_{t+h,i},\hat{\gamma}_{t+h,i})\\ =\frac{1}{N}\sum_{i=1}^{N}|\frac{d\hat{I}_{t+h,i}}{d}-(\hat{\beta}_{t+h,i}\hat{S}_{t+h,i}\hat{I}_{t+h,i}-\hat{\gamma}_{t+h,i}\hat{I}_{t+h,i})|,\\ \mathcal{L}_{p}=\mathcal{L}_{d}(\hat{I}_{t+h,i})+\mathcal{L}_{o}(\hat{S}_{t+h,i},\hat{I}_{t+h,i},\hat{\beta}_{t+h,i},\hat{\gamma}_{t+h,i}),\end{gathered}$ (11)
where $x_{t+h,i}$ is the ground truth. $\mathcal{L}_{d}$ represents the error between predicted values and ground truth, which uses mean absolute error (MAE) as a loss measure. $\mathcal{L}_{o}$ is the residual term of the ODEs set. We use $\mathcal{L}_{d}$ to constrain the output of MLPs consistent with the real-world epidemic transmission dynamics and use $\mathcal{L}_{o}$ to constrain the output consistent with the mechanism of intrinsic epidemic transmission.
In our framework, we use the forecasting from the epidemic forecasting decoder module as the final forecasting result and use the MAE loss to compare it with the ground truth.
$\mathcal{L}_{g}=\frac{1}{N}\sum_{i=1}^{N}|\hat{y}_{t+h,i}-x_{t+h,i}|,$ (12)
where $\hat{y}_{t+h,i}\in\hat{\boldsymbol{Y}}_{t+h}$ is output value of HeatGNN and $x_{t+h,i}$ is the ground truth value. Finally, the main objective function $\mathcal{L}$ of HeatGNN is given as follows:
$\mathcal{L}=\mathcal{L}_{g}+\lambda\mathcal{L}_{p},$ (13)
where the loss term $\mathcal{L}_{g}$ is forecasting the error of our proposed framework and minimizing it to achieve good performance. Another loss term $\mathcal{L}_{p}$ is used as a regularizer of the EIEL module to enhance the ability of this module to represent the intrinsic epidemic transmission mechanism. The loss weight $\lambda$ is used to balance two different loss terms.
IV Experiments
In the experiment, we evaluate our proposed framework HeatGNN by answering the below research questions (RQs):
  • RQ1: Does HeatGNN outperform the existing state-of-the-art epidemic forecasting models?
  • RQ2: What is the impact of each core component in HeatGNN?
  • RQ3: What is the performance of HeatGNN with different hyperparameters?
  • RQ4: Is HeatGNN reliable for short-term and long-term epidemic forecasting?
  • RQ5: Does HeatGNN have robust scalability?
IV-A Datasets
We evaluate the performance of our framework using three real-world datasets following previous work[4]: Japan-Prefectures, US-States, and US-Regions.
  • Japan-Prefectures: This dataset comprises weekly influenza-like-illness levels from August 2012 to March 2019 for 47 prefectures in Japan collected from the Weekly Infectious Diseases Report in Japan.
  • US-Regions: This dataset includes weekly influenza activity levels (flu patient counts) from 2002 to 2017 for 10 Health and Human Services regions in the United States collected from the Department of Health and Human Services.
  • US-States: This dataset contains weekly influenza disease data (positive cases) from 2010 to 2017 for 49 states in the United States collected from the Center for Disease Control.
The statistics of these three datasets are summarized in Table I. The US-Regions dataset has a higher mean and greater data volatility, the US-States dataset has a lower mean and more concentrated data distribution, and the Japan-Prefectures dataset is between the other two regarding numerical range and data dispersion.
TABLE I: Statistics of the three real-world datasets.
Dataset Size Min Max Mean SD
Japan-Prefectures $47\times 348$ 0 26635 655 1711
US-Regions $10\times 785$ 0 16526 1009 1351
US-States $49\times 360$ 0 9716 223 428
IV-B Evaluation Metrics
The Root Mean Squared Error (RMSE) measures the difference between predicted and ground truth values after projecting the normalized values into the real range:
$\mathrm{RMSE}=\sqrt{\frac{1}{N}\sum_{i=1}^{N}\left(\hat{y}_{i}-y_{i}\right)^{2}},$ (14)
The PCC is a measure of the linear dependence between two sets of values:
$\mathrm{PCC}=\frac{\sum_{i=1}^{N}\left(\hat{y}_{i}-\overline{\hat{y}}\right)\left(y_{i}-\bar{y}\right)}{\sqrt{\sum_{i=1}^{N}\left(\hat{y}_{i}-\overline{\hat{y}}\right)^{2}}\sqrt{\sum_{i=1}^{N}\left(y_{i}-\bar{y}\right)^{2}}}.$ (15)
where $\hat{y}$ is the predicted values, $y$ is the ground truth values, and $\overline{\hat{y}}$ and $\bar{y}$ are their average values respectively. $N$ represents the number of samples.
IV-C Experimental Settings and Implementation Details
We divide the dataset into training, validation, and test sets in chronological order with a ratio of 60%-20%-20%[29, 20]. We normalized all data based on the training data. The validation set is used for hyperparameter tuning and to prevent overfitting. The test set is used to evaluate the performance of the final model using RMSE and PCC[29]. The smaller the RMSE, the better the performance while the larger the PCC, the better the performance. We used Adam [30] as the optimizer with weight decay 5e-4 during training. The grid search is used to find the best hyperparameter combinations. After the grid search, we set the initial learning rate to 0.001, the early stopping strategy with the patience of 200 epochs, the batch size to 32, and the number of epochs to 1500. All the parameters are initialized using Xavier initialization in HeatGNN. For all models, we set the historical window size $w$ to 20, that is, to use the historical data of the previous 20 weeks as input features. The horizon of advance forecasting $h$ is 2, 3, 4, and 5, that is, to predict the risk of epidemic transmission of the $h$-th week after time step $t$. All experimental results are the average of 5 randomized trials. The experiments are conducted on an Intel Xeon Silver 4214 Processor CPU with 128 GB of 2666 MHz RAM and 4 NVIDIA TITAN RTX 24GB GPUs. Our method implemented in Python 3.8 and PyTorch 1.13.1 trains on GPUs in 30-60 minutes for each forecasting task.
IV-D Comparison Methods
We compared four types of models: traditional statistical models, RNN-based and CNN-based models, STGNN models, and EML models. The traditional statistical models include AutoRegressive (AR), AutoRegressive Moving Average (ARMA), Generalized AutoRegressive (GAR), and Vector AutoRegression (VAR). The RNN-based and CNN-based models include Recurrent Neural Network (RNN)[31], Long Short-Term Memory (LSTM)[32], Gated Recurrent Unit (GRU)[33], RNN-Attn[34], CNNRNN-Res[29], and Long- and Short-term Time-series Network (LSTNet)[35]. The STGNN models include STGCN[27], Multiresolution Graph Neural Networks (MGNN)[15], Temporal Multiresolution Graph Neural Networks (TMGNN)[15], Attention-based Multiresolution Graph Neural Networks (ATMGNN)[16], Cross-location Attention-based Graph Neural Networks (Cola-GNN)[4], and Epi-GNN[12]. The EML model is Epi-Cola-GNN[20].
TABLE II: RMSE and PCC performance of different methods on three datasets with horizon = 2, 3, 4, 5. A boldface indicates the best result for each column, and the second-best is underlined.
Japan-Prefectures US-Regions US-States
RMSE($\downarrow$) 2 3 4 5 2 3 4 5 2 3 4 5
AR 1669.898 2071.535 2312.712 2440.119 625.620 813.453 956.083 1078.889 172.179 210.689 239.205 264.437
ARMA 1661.222 2070.781 2312.378 2438.737 622.393 804.126 939.212 1057.998 172.106 207.789 239.828 262.015
GAR 1484.372 1990.001 2281.875 2427.555 607.232 790.684 934.327 1053.783 161.848 199.566 220.901 247.398
VAR 1577.216 2100.213 2443.453 2566.043 694.516 864.829 1000.032 1112.857 297.669 306.723 318.044 333.312
RNN 1198.314 1551.096 1720.115 1851.500 572.438 753.568 890.801 979.708 158.949 195.903 210.525 224.040
LSTM 1255.470 1555.815 1710.241 1839.544 570.277 743.945 913.847 1048.754 161.029 197.831 216.700 226.667
GRU 1181.230 1498.437 1632.849 1781.807 570.898 741.786 893.440 1024.078 159.298 195.385 217.282 247.858
RNN-Attn 1311.157 1751.819 2056.583 2266.540 566.375 757.980 894.176 999.389 162.821 193.795 219.705 243.932
CNNRNN-Res 1401.394 1889.815 2260.501 2570.203 594.910 782.389 873.559 1006.251 228.275 290.916 308.268 320.370
LSTNet 1441.367 2030.374 2331.407 2385.490 606.540 841.441 1030.450 1134.228 195.949 232.584 255.764 276.306
STGCN 1273.563 1389.538 1498.233 1399.753 726.945 814.875 894.509 965.067 205.060 226.799 246.035 260.379
MGNN 1756.653 1922.766 1822.108 1943.240 1472.420 1415.071 1473.778 1487.338 178.034 189.710 198.840 205.517
TMGNN 1500.051 1471.865 1568.222 1558.965 656.364 755.206 829.295 915.758 178.709 188.653 201.361 190.089
ATMGNN 2436.255 2465.356 2473.451 2464.829 1553.142 1607.391 1691.877 1688.661 307.932 374.719 383.963 355.191
Cola-GNN 1167.715 1416.884 1478.782 1572.566 551.810 681.407 793.208 870.818 147.700 175.602 194.336 223.540
Epi-Cola-GNN 1118.387 2056.956 2282.738 1543.628 546.637 747.352 881.656 871.939 148.455 184.681 204.389 219.132
Epi-GNN 1503.244 1296.059 1284.933 1447.863 563.328 708.990 790.972 881.026 149.515 168.939 184.500 190.005
HeatGNN 1148.850 1260.150 1295.697 1378.207 540.573 693.655 781.513 852.311 142.433 164.788 177.624 186.280
PCC($\uparrow$) 2 3 4 5 2 3 4 5 2 3 4 5
AR 0.743 0.567 0.413 0.303 0.921 0.868 0.829 0.788 0.934 0.903 0.876 0.850
ARMA 0.745 0.567 0.413 0.303 0.920 0.870 0.827 0.786 0.933 0.902 0.873 0.849
GAR 0.803 0.613 0.444 0.318 0.924 0.870 0.823 0.782 0.938 0.906 0.884 0.861
VAR 0.771 0.564 0.369 0.229 0.895 0.839 0.775 0.715 0.779 0.769 0.749 0.730
RNN 0.887 0.810 0.775 0.757 0.934 0.886 0.848 0.819 0.944 0.915 0.902 0.890
LSTM 0.884 0.831 0.818 0.820 0.933 0.890 0.853 0.820 0.943 0.915 0.897 0.892
GRU 0.894 0.840 0.827 0.807 0.933 0.887 0.849 0.817 0.944 0.916 0.897 0.870
RNN-Attn 0.865 0.735 0.589 0.497 0.936 0.889 0.849 0.820 0.943 0.916 0.891 0.869
CNNRNN-Res 0.828 0.653 0.459 0.239 0.925 0.869 0.833 0.777 0.887 0.802 0.769 0.731
LSTNet 0.820 0.586 0.384 0.338 0.934 0.877 0.795 0.733 0.924 0.885 0.848 0.811
STGCN 0.892 0.874 0.852 0.872 0.899 0.866 0.848 0.831 0.901 0.882 0.861 0.838
MGNN 0.688 0.633 0.644 0.507 0.558 0.557 0.557 0.550 0.926 0.914 0.909 0.901
TMGNN 0.827 0.836 0.833 0.822 0.913 0.885 0.870 0.833 0.925 0.916 0.906 0.913
ATMGNN 0.160 0.156 0.162 0.170 0.496 0.481 0.464 0.456 0.737 0.640 0.633 0.698
Cola-GNN 0.906 0.845 0.870 0.834 0.939 0.903 0.872 0.847 0.953 0.936 0.922 0.904
Epi-Cola-GNN 0.912 0.580 0.444 0.859 0.940 0.889 0.851 0.858 0.953 0.922 0.906 0.909
Epi-GNN 0.821 0.911 0.911 0.869 0.937 0.904 0.885 0.858 0.952 0.937 0.925 0.923
HeatGNN 0.917 0.907 0.910 0.884 0.941 0.908 0.885 0.866 0.953 0.939 0.929 0.921
IV-E Overall Comparison
This section addresses RQ1. We evaluate our method in both short-term forecasting (horizon=2, 3) and long-term forecasting (horizon=4, 5) settings. We exclude the case of horizon=1 due to delays in epidemic surveillance data, as extreme short-term forecasting does not provide sufficient time for effective interventions. Table II summarizes the results of all methods under three real-world datasets. While the PCC shows relatively stable ranges and variances between datasets, the RMSE shows significant variability, primarily driven by differences in the variance of the datasets.
Compared with various baseline models, including traditional statistical, RNN-based, CNN-based, STGNN, and EML models, HeatGNN achieves the lowest RMSE and highest PCC across most horizons and datasets. These results demonstrate its capability to capture both spatio-temporal heterogeneity and mechanistic heterogeneity effectively. The high PCC values across datasets and horizons highlight the robustness of HeatGNN in capturing transmission patterns that are closely related to real-world data, making it well-suited for short-term and long-term epidemic forecasting. For instance, in the Japan-Prefectures dataset, HeatGNN achieves a PCC of 0.917 for horizon=2, outperforming the next best model, Epi-Cola-GNN, which scores 0.902. Among the other models, Epi-Cola-GNN and Epi-GNN show the second-best performances. While the forecasting accuracy of Epi-GNN is slightly inferior to HeatGNN in most tasks, it achieves the best performance in a few cases, highlighting its effectiveness and robustness in capturing spatio-temporal dependencies for epidemic forecasting.
As the forecasting horizon increases, HeatGNN maintains competitive performance, while several baseline models show significant error growth. Short-term forecasting performs well due to strong temporal dependencies, as evidenced by good forecasting performance for all models at horizon=2. However, traditional statistical and CNN-based models lose temporal dependency information as the horizon extends, leading to rapid performance declines. In contrast, the RNN-based models maintain relatively robust performance with their memory of temporal dependencies. The improved performance of STGNN models, such as Cola-GNN, STGCN, and TMGNN, highlights the importance of spatio-temporal heterogeneity in epidemic forecasting. Nevertheless, while spatio-temporal dependencies enhance the ability to capture dynamic changes in epidemic transmission, they also introduce added complexity, compromising the robustness of specific models, such as MGNN and ATMGNN. Additionally, incorporating epidemic mechanism models into GNNs improves short-term forecasting accuracy. For instance, Epi-Cola-GNN achieves the second-best performance across all three datasets at horizon=2.
IV-F Ablation Study
This section addresses RQ2. To assess the effect of each component in our framework, we conduct the ablation study by testing the following variants of HeatGNN:
  • HeatGNN w/o PL: Remove the physics loss from HeatGNN removes the constraints on the EIEL module.
  • HeatGNN w/o TG: Remove the TG module and use only the STGL module for direct forecasting, while retaining the EIEL module to guide the output of the framework.
  • HeatGNN w/o TP: Remove the EIEL and TG modules and only keep the STGL module for forecasting.
The ablation study results of RMSE and PCC are shown in Table III. The results show that our proposed HeatGNN performs robustly across most cases, except for horizon=4 in the US-States dataset. The physical loss incorporates domain knowledge of epidemic transmission, particularly benefiting high-variability regions (e.g., Japan-Prefectures and US-Regions) and long-term forecasting. Removing the physics loss results in the model lacking intrinsic epidemic transmission mechanism constraints, which increases the forecasting error and reduces its ability to capture trends. However, its impact is limited in the US-States dataset, which is low volatility and relatively uniform data. The TG module is crucial for modeling cross-regional epidemic transmission. After removing the TG module, the model is significantly weaker at capturing the intrinsic epidemic transmission relationship between locations. While it has a similar role to the physics loss across datasets, its impact on forecasting performance is more pronounced. Additionally, the EIEL module enhances the effectiveness of the TG module, further improving the model’s ability to capture epidemic transmission. Relying solely on the STGL module is insufficient for addressing complex epidemic forecasting tasks. When both the TG and EIEL modules are removed, the model struggles to accurately capture epidemic transmission over extended time horizons, resulting in a significant decline in forecasting performance. These findings suggest that the modules interact synergistically to achieve optimal performance in HeatGNN.
TABLE III: Ablation test results on three datasets.
RMSE($\downarrow$) 2 3 4 5
Japan-Prefectures
HeatGNN w/o PL 1230.134 1360.714 1407.590 1453.051
HeatGNN w/o TG 1228.951 1331.692 1421.312 1438.525
HeatGNN w/o TP 1278.808 1322.376 1441.331 1451.568
HeatGNN 1148.850 1260.150 1295.697 1378.207
US-Regions
HeatGNN w/o PL 631.080 799.733 869.248 942.947
HeatGNN w/o TG 739.933 811.149 854.826 1039.151
HeatGNN w/o TP 726.639 861.203 947.868 977.450
HeatGNN 540.573 693.655 781.513 852.311
US-States
HeatGNN w/o PL 146.414 168.938 172.590 188.129
HeatGNN w/o TG 148.601 176.770 176.644 201.611
HeatGNN w/o TP 152.508 181.273 169.937 206.313
HeatGNN 142.433 164.788 177.624 186.280
PCC($\uparrow$) 2 3 4 5
Japan-Prefectures
HeatGNN w/o PL 0.900 0.889 0.876 0.863
HeatGNN w/o TG 0.891 0.895 0.898 0.871
HeatGNN w/o TP 0.890 0.893 0.893 0.853
HeatGNN 0.917 0.907 0.910 0.884
US-Regions
HeatGNN w/o PL 0.924 0.892 0.875 0.864
HeatGNN w/o TG 0.923 0.891 0.881 0.838
HeatGNN w/o TP 0.927 0.895 0.866 0.854
HeatGNN 0.941 0.908 0.885 0.866
US-States
HeatGNN w/o PL 0.951 0.933 0.929 0.917
HeatGNN w/o TG 0.951 0.927 0.930 0.907
HeatGNN w/o TP 0.951 0.925 0.932 0.902
HeatGNN 0.953 0.939 0.929 0.921
IV-G Hyperparameter Sensitivity Analysis
This section addresses RQ3. We investigate how the forecasting performance varies with some hyperparameters.
IV-G1 Window size
To evaluate whether the model is sensitive to the length of historical data, we evaluate different window sizes ranging from 20 to 60. Figures 3 (a), (b), and (c) show the RMSE and PCC results in different window sizes. For the Japan-Prefectures dataset, forecasting performance improves significantly with larger window sizes, as evidenced by a gradual decrease in RMSE and an increase in PCC. This result suggests that a larger window size provides more comprehensive historical information, enhancing forecasting accuracy. In contrast, model performance remains stable across all window sizes for the US-Regions dataset, which indicates that the model is not sensitive to the window size change in this dataset. For the US-States dataset, the optimal performance is observed at a window size of 60. In summary, while increasing the window size enhances model performance, the improvement becomes limited after a threshold.
IV-G2 Horizon size
To evaluate the sensitivity of the model to the size of the forecasting horizon, we analyze its performance across horizons ranging from 1 to 6. Figures 3 (d), (e), and (f) show the RMSE and PCC results for different horizons. In the Japan-Prefectures dataset, as the forecasting horizon increases, RMSE gradually rises while PCC decreases, indicating a decline in forecasting accuracy with longer horizons. Similar trends are observed in the US-Regions and US-States datasets. Comparing the results across the three datasets, we find that datasets with more significant fluctuations (Japan-Prefectures), shorter forecasting horizons, and more accurate predictions, while datasets with lower variability (US-Regions and US-States) maintain relatively robust forecasting performance even with longer horizons.
IV-G3 Loss weight
To evaluate the sensitivity of the model to the balance weight between physical and data loss, we examine the effect of varying $\lambda$ values from 0 to 1 on forecasting performance. Figures 3 (g), (h), and (i) show the RMSE and PCC results under different $\lambda$. On the Japanese-Prefectures dataset, the PCC of the model remains stable as $\lambda$ increases (i.e., the weight of the physical loss increases), while the RMSE increases slightly at $\lambda$= 0.2. This result indicates that performance is slightly affected by low physical loss weights, and the overall trend remains stable. Comparing the results across the three datasets, we observe that the Japan-Prefectures dataset performs slightly worse under low physical loss weights. In contrast, the US-States dataset performs best with higher physical loss weights. Overall, increasing the $\lambda$ value has a greater impact on datasets with greater variability, while datasets with lower variability (e.g., US-States) maintain good performance even at higher $\lambda$ values. These results suggest that reasonably adjusting the balance between physical and data loss can enhance model performance across different datasets.
IV-G4 Sparse Threshold
To evaluate whether the model is sensitive to the sparsification of the MAG, we test the impact of different sparse thresholds, from 0 to 0.8 (Figures 3 (j), (k), and (l)). A higher threshold corresponds to a sparser MAG. On the Japan-Prefectures dataset, as the sparse threshold increases, RMSE fluctuates slightly, indicating that excessive sparsification may degrade the forecasting performance. On the US-Regions dataset, RMSE shows minor fluctuation, with mild sparsification improving model performance, while PCC remains stable, indicating that the model is not highly sensitive to changes in the sparse threshold. For the US-States dataset, the model achieves the best at a lower sparse threshold, with a slight decrease in PCC as the threshold increases. In summary, a modest increase in the sparse threshold can improve model performance, and selecting an appropriate threshold helps balance data complexity and model effectiveness.
(a) Japan-Prefectures
(b) US-Regions
(c) US-States
(d) Japan-Prefectures
(e) US-Regions
(f) US-States
(g) Japan-Prefectures
(h) US-Regions
(i) US-States
(j) Japan-Prefectures
(k) US-Regions
(l) US-States
Figure 3: Sensitivity analysis on hyperparameter in three real-world datasets. Performance with window size $w$, horizon size $h$, loss weight $\lambda$, and sparse threshold $\delta$. The shaded area represents the 95% confidence interval. Except for the sensitivity analysis experiment of $h$, other sensitivity analysis experiments are conducted under horizon=5.
IV-H Case Study
This section addresses RQ4. To demonstrate the effectiveness of the proposed HeatGNN in capturing spatio-temporal heterogeneity and mechanistic heterogeneity, we conduct a case study to evaluate its short-term and long-term forecasting performance. Specifically, we select two locations from the Japan-Prefectures dataset. We analyze their forecasting performance under the short-term (horizon=2) and long-term (horizon=5) scenarios to explore the forecasting performance of the HeatGNN under different locations. In short-term forecasting (Figures 4(a) and 4(b)), the model accurately predicts the fluctuation trends of epidemic transmission risk. The model’s predictions are close to the actual situation both at the peak and trough of the epidemic, indicating that it is also applicable during extreme events (e.g., 8-10 weeks in Figure 4(a)). This peak sensitivity is critical in the epidemic response strategy, which helps to take interventions in advance to control outbreaks. The model performs stable prediction performance in long-term forecasting, in which the model captures the direction of risk fluctuations (Figures 4(c) and 4(d)). Furthermore, the model can predict a rapid rise in the epidemic 1-2 weeks before an outbreak in long-term forecasting, providing valuable lead time for early prevention efforts. To further validate our findings, we perform a similar analysis using data from the US-States dataset (Figure 5), and the results are consistent with those from the Japan-Prefectures dataset. Those results indicate that HeatGNN effectively captures both short-term and long-term epidemic transmission.
(a) No.0 Location
(b) No.25 Location
(c) No.0 Location
(d) No.25 Location
Figure 4: Case study in Japan-Prefectures. (a) and (b) are conducted under horizon=2. (c) and (d) are conducted under horizon=5.
(a) No.0 Location
(b) No.7 Location
(c) No.0 Location
(d) No.7 Location
Figure 5: Case study in US-Regions. (a) and (b) are conducted under horizon=2. (c) and (d) are conducted under horizon=5.
IV-I Inference Analysis
This section addresses RQ5. To verify the inference efficiency of our framework, we conduct an inference analysis experiment of different location scales based on the US-States dataset to evaluate the inference speed. The results in Figure 6 indicate that the inference time of the HeatGNN remains stable with only minor fluctuations under different location scales. This consistent performance demonstrates the robustness and reliability of HeatGNN, further highlighting its scalability and suitability for applications involving dynamic or large-scale epidemic transmission graph datasets.
Figure 6: Inference Analysis with different scale locations under horizon=5 in US-Regions.
V Related Work
V-A Epidemic Forecasting
Epidemic forecasting is a critical task in monitoring the spread of epidemics to help public health interventions quickly and avoid large-scale outbreaks [36]. The time dependence and seasonality of epidemic spread are considered at the beginning of epidemic forecasting. Time series regression models, such as ARIMA[37] and VAR [38], and machine learning models, such as random forests[5], are widely used to solve this task, but these models cannot capture spatial dependencies and ignore the evolution of epidemic transmission. To further improve the ability of the model to capture temporal dependencies, the RNN and RNN-based models were proposed for this task [39]. CNN, Transformer, and their adopted models are used to capture multi-scale features, such as short-term temporal dependencies and long-term temporal dependencies[40]. Meanwhile, the emergence of GNN provides an effective solution to identifying spatio-temporal patterns in epidemic forecasting. In addition to most studies focusing on the impact of geographic topology between locations [4, 14, 12], some studies focus on the distinction between local and global information in epidemic spread [12], as well as multi-scale information [13, 41]. However, these methods mainly focus on the spatio-temporal dependence of epidemic transmission but ignore intrinsic transmission patterns and the evolution of epidemics. They also lack mechanism models to describe the transmission mechanisms behind epidemics. These limitations make it challenging to provide sufficient information to support public health decision-making.
V-B Spatio-temporal Graph Learning
Recently, spatio-temporal graph learning has emerged as a powerful framework for modeling complex data structures that evolve over both spatial and temporal. Graph Convolutional Networks (GCNs) have been developed to achieve end-to-end learning of graph structures[42]. GCNs can efficiently capture local features by sharing parameters in the neighborhood of the graph. Since then, GNNs expanded from static graphs to dynamic graphs to handle the requirement of spatio-temporal data modeling from the needs of dynamic network tasks, like traffic[43], electric load[44], and epidemic[4] forecasting, which involves temporal evolution. The emergence of dynamic graph neural networks, like graph convolutional recurrent networks (GCRN), enables GNNs to update graph structures in time series and capture dynamic changes in node relationships[45]. To further integrate temporal and spatial dependencies, STGNNs combine graph convolution and temporal convolution, enabling the model to capture the dynamic characteristics of temporal and spatial simultaneously, showing significant advantages in various spatio-temporal data tasks [27].
With the development of spatio-temporal graph learning, some studies have introduced recurrent neural networks [46, 47] and Transformer-based attention mechanisms [48] to learn spatio-temporal dependencies more flexibly. In addition, the demand for spatio-temporal dependencies across different scales is driven by the proposed multiscale STGNNs[49]. STGNNs combined with the hypergraph structure are used to process complex high-order interactions in spatio-temporal graph learning[50, 51]. The spatio-temporal graph diffusion network was proposed to describe the information diffusion process in spatio-temporal graph learning [43]. Meanwhile, he heterogeneous STGNNs have been proposed to better adapt to complex heterogeneity problems in spatio-temporal data[52]. Overall, STGNN can capture the spatio-temporal dependencies of epidemics well. Therefore, to enable our proposed framework to capture the spatio-temporal heterogeneity of various epidemic locations, we are inspired by [12] and use it as the spatio-temporal graph module.
V-C Physics-guided Machine Learning
The approximations learned by machine learning methods often ignore or violate existing physical knowledge, leading to misleading or unexplainable conclusions [53]. Therefore, the combination of physics and machine learning has begun to be proposed to solve the problem of lack of interpretability and black-box structures in machine learning. A commonly used method is physics-informed neural networks (PINN) [28], which is based on combining physical residual loss with data loss to transfer or distill physical epidemic knowledge into a machine learning framework for constrained learning processes while being able to effectively capture heterogeneous data in different forecasting tasks, such as urban flow[54], blood pressure[55], rumor detection[56], and epidemic transmission[57, 19]. Another way to combine physical models and machine learning is Neural differential equations (Neural ODEs), which use the neural network as a continuous dynamic process. For example, the time-dependent states in the epidemic model are replaced by neural networks[58]. Neural ODEs are also used in fields such as climate, weather[59], polar motion[60], and battery health[61] forecasting to solve real-world complex continuous-time systems.
Some scenarios in epidemic transmission require integrating physical models into machine learning models to improve performance and interpretability. Epidemic models are based on overly simple assumptions and cannot fully understand the complex epidemic spread[9], while machine learning models lack effective mechanism explanations. Extended epidemic models with networks have been proposed to describe the transmission patterns between locations[62, 11]. However, these network-based epidemic models have problems with large-scale parameter optimization[11] or lack of data[63, 21]. To address those problems, some models proposed to use simulated data generated by physical models to solve the problem of insufficient data and provide high-resolution forecasting[17, 63]. Meanwhile, machine learning has also begun to estimate epidemic models to solve the parameter explosion [21, 20]. Neural ODEs are incorporated into epidemic models to learn the continuous evolution of epidemic spread in different regions [18]. The epidemic models were used as causal mechanism models to guide the graph embedding learning process to handle noisy data and missing data better[64]. Although physics-guided machine learning has made efforts in epidemic forecasting, previous studies have ignored the mutual influence and dependency of the intrinsic epidemic transmission between locations, which makes it challenging to forecast epidemics breaking out in multiple locations.
VI Conclusion
In this study, we propose a novel epidemic forecasting framework, HeatGNN, to improve the forecasting performance. We incorporate the epidemic mechanism model into the epidemiology-informed embedding learning to learn intrinsic epidemic transmission patterns. We use the physics loss to constrain the epidemiology-informed location embeddings to align with real-world epidemics. We propose a mechanistic affinity graph to represent the mechanistic dependence of intrinsic epidemic transmission between locations. We design a heterogeneous transmission graph network to account for the mechanistic heterogeneity. Finally, the forecasting performance is further improved by combining the spatio-temporal graph learning and transmission graph modules. The effectiveness of HeatGNN is validated using real-world datasets for epidemic forecasting. Experimental results demonstrate that HeatGNN is capable of outperforming existing state-of-the-art models in epidemic forecasting. In the future, we plan to construct an interpretable mechanistic affinity graph with mechanisms models to cope with the spatio-temporal and mechanistic heterogeneity.